OpenStudy (aaronandyson):

Prove that secA(1-sinA)(secA+tanA)=1?

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (michele_laino):

here we have to use these identities: \[\Large \sec A = \frac{1}{{\cos A}},\quad \tan A = \frac{{\sin A}}{{\cos A}}\]

OpenStudy (michele_laino):

hint: we can write this step: \[\Large \begin{gathered} \sec A\left( {1 - \sin A} \right)\left( {\sec A + \tan A} \right) = \hfill \\ \hfill \\ = \frac{1}{{\cos A}}\left( {1 - \sin A} \right)\left( {\frac{1}{{\cos A}} + \frac{{\sin A}}{{\cos A}}} \right) = ...? \hfill \\ \end{gathered} \] please, simplify

OpenStudy (aaronandyson):

???

OpenStudy (michele_laino):

I have replaced secA wit 1/cosA and tanA with (sinA)/(cosA)

OpenStudy (aaronandyson):

ok!

OpenStudy (aaronandyson):

what next?>.<

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} \sec A\left( {1 - \sin A} \right)\left( {\sec A + \tan A} \right) = \hfill \\ \hfill \\ = \frac{1}{{\cos A}}\left( {1 - \sin A} \right)\left( {\frac{1}{{\cos A}} + \frac{{\sin A}}{{\cos A}}} \right) = \hfill \\ \hfill \\ = \frac{{1 - \sin A}}{{\cos A}} \cdot \frac{{1 + \sin A}}{{\cos A}} = ...? \hfill \\ \end{gathered} \] now it is easy to write the next step

OpenStudy (aaronandyson):

cross multiply?

OpenStudy (michele_laino):

you have to multiply those two fractions

OpenStudy (aaronandyson):

??

OpenStudy (michele_laino):

hint: \[\large \begin{gathered} \sec A\left( {1 - \sin A} \right)\left( {\sec A + \tan A} \right) = \hfill \\ \hfill \\ = \frac{1}{{\cos A}}\left( {1 - \sin A} \right)\left( {\frac{1}{{\cos A}} + \frac{{\sin A}}{{\cos A}}} \right) = \hfill \\ \hfill \\ = \frac{{1 - \sin A}}{{\cos A}} \cdot \frac{{1 + \sin A}}{{\cos A}} = \hfill \\ \hfill \\ = \frac{{\left( {1 - \sin A} \right)\left( {1 + \sin A} \right)}}{{{{\left( {\cos A} \right)}^2}}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (aaronandyson):

1-sin^2/cos^2 is 1

OpenStudy (michele_laino):

that's right!

OpenStudy (aaronandyson):

thanks help me more?

OpenStudy (michele_laino):

ok!